OpenStudy (anonymous):

8^(x+8) =32^x Can someone tell me how I might be able to solve this?

OpenStudy (anonymous):

\[8^{x+8}=32^{x}\]

OpenStudy (anonymous):

Thats what it looks like

OpenStudy (lgbasallote):

(2^3)^(x+8) = 2^5^(x) 3(x+8) = 5x 3x + 24 = 5x 24 = 2x x= 12 do you get the process?

OpenStudy (anonymous):